# LRT-Diffusion: Calibrated Risk-Aware Guidance for Diffusion Policies

### v1

@@ Appendix B Proofs for Finite-sample Calibration and Stability
B.1 Finite-sample Calibration: Proof of Theorem 2
@@
B.2 Deterministic Displacement bound: Proof for Lemma 3
@@
B.3 One-step LLR Decomposition and Moment Bounds
@@
$\Delta\ell_{t}=\underbrace{\frac{\Delta\mu_{t}^{\top}\!\left(m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\right)}{\sigma_{t}^{2}}}_{=:~A_{t}\text{ (deterministic given }\mathcal{F}_{t})}\;+\;\underbrace{\frac{\Delta\mu_{t}^{\top}z_{t}}{\sigma_{t}}}_{=:~B_{t}\text{ (zero-mean Gaussian)}}.$
@@
where in the last inequality we used the displacement bound $\|m_{t}-\mu_{u,t}\|\leq B_{t}$ from §B.2.
@@
B.4 Sub-Gaussian concentration for the cumulative LLR
@@
Under the scale controls in §B.2,
@@
The cap $\beta_{\max}$ and gradient clipping (via $G,\lambda_{\max}$) yield a per-step deterministic displacement $B_{t}$ and a uniform bound $B_{\text{step}}$; together with the equal-covariance identity, these give (i) explicit bounds on the conditional mean and variance of each LLR increment, and (ii) a sub-Gaussian concentration for the cumulative evidence with proxy $V=\sum_{t}\|\Delta\mu_{t}\|^{2}/\sigma_{t}^{2}$. These stability controls make the calibration reproducible (Theorem 2) and the gate numerically well-conditioned.
Appendix C When LRT Dominate Q guidance
C.1 Proof for Proposition  3
@@
C.2 Condition for LRT to Dominate
Corollary 3.1 (Feasible $\alpha$ via the union bound).
Under Assumption 4 with $c(\alpha)\leq\alpha T$, a sufficient condition for $\Delta_{\text{true}}\geq 0$ is
$\alpha~\leq~\alpha_{\max}:=\frac{\Delta_{\widehat{Q}}-2\varepsilon_{\rm in}-\Delta_{\varepsilon}\,\eta(\pi_{Q})}{\Delta_{\varepsilon}\,T}.$ (24)
@@ C.2 Condition for LRT to Dominate
Plug $\eta(\pi_{\rm LRT})\leq\alpha T$ into (23) and solve $\Delta_{\text{true}}\geq 0$ for $\alpha$. Since $\Delta_{\varepsilon}T>0$ unless $T=0$ or $\Delta_{\varepsilon}=0$, (24) follows. ∎
@@
C.3 A tighter, Data-driven Alternative
The surrogate $\eta(\pi_{\rm LRT})\leq\alpha T$ is often loose. Calibrate once and measure $\widehat{\eta}_{\rm LRT}(\alpha)$ on held-out states (our kNN proxy). Then the same derivation yields the data-driven bound
$\Delta_{\text{true}}~\geq~\Delta_{\widehat{Q}}~-~2\varepsilon_{\rm in}~-~\Delta_{\varepsilon}\big(\widehat{\eta}_{\rm LRT}(\alpha)+\widehat{\eta}(\pi_{Q})\big),$ (25)
@@ C.3 A tighter, Data-driven Alternative
C.4 Optional Refinement with Consistency or Distance-based Bounds
@@
$\Delta_{\text{true}}~\geq~\Delta_{\widehat{Q}}~-~2r_{n}~-~\Delta_{\varepsilon}\big(\eta(\pi_{\rm LRT})+\eta(\pi_{Q})\big),$
@@
$\Delta_{\text{true}}~\geq~\Delta_{\widehat{Q}}-2\varepsilon_{\rm in}-L_{Q}\,\mathbb{E}_{s}\!\big[\mathrm{dist}(a_{\rm LRT},\mathcal{S}(s))+\mathrm{dist}(a_{Q},\mathcal{S}(s))\big].$
@@
C.5 Further Discussion for LRT dominate Q
Combining Prop. 3 and Prop. 4 yields the sufficient condition
$\Delta_{Q^{\rm true}}~\geq~\Delta_{\hat{Q}}~-~2\varepsilon_{\rm in}~-~\kappa\big(\alpha T+\eta(\pi_{Q})\big),$
$\alpha_{\max}~=~\frac{\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\kappa\,\eta(\pi_{Q})}{\kappa\,T},$
@@ C.5 Further Discussion for LRT dominate Q
  • OOD penalty $\kappa(\alpha T+\eta(\pi_{Q}))$. The LRT part scales with $\alpha$ and with the number of gated steps $T$ via the union proxy $\eta(\pi_{\rm LRT})\leq\alpha T$; the Q–guided baseline pays $\kappa\,\eta(\pi_{Q})$. Smaller $\kappa$—e.g., via a smaller $\beta_{\max}$ or late–step gating when $\sigma_{t}$ is small—makes the condition easier to satisfy.
C.5.1 When does LRT “dominate” Q–guidance?
@@
  • A small–OOD Q baseline helps the certificate. If $\eta(\pi_{Q})$ is small (Q–guidance rarely leaves support), the denominator in $\alpha_{\max}$ is fixed while the numerator increases, so $\alpha_{\max}$ grows: $\eta(\pi_{Q})\!\downarrow\Rightarrow\alpha_{\max}\!\uparrow$. In the extreme $\eta(\pi_{Q}){=}0$, the feasible range reads $\alpha\leq(\Delta_{\hat{Q}}-2\varepsilon_{\rm in})/(\kappa T)$. Intuitively, when Q–guidance already behaves conservatively, we can open the LRT gate more (choose a larger $\alpha$) and still retain a certificate.
  • Levers that enlarge $\alpha_{\max}$. (i) Reduce $T$ by gating only in late steps (then $\alpha_{\max}\propto 1/T$); (ii) Lower $\kappa$ (e.g., smaller $\beta_{\max}$ or restricting gating to small–variance steps; anchoring gradients at $\mu_{u}$ also lowers off–support effects); (iii) Increase $\Delta_{\hat{Q}}$ by taking a modest Q–step near $\mu_{\rm LRT}$ when the gate opens or by using evidence–tied/annealed guidance.
C.5.2 Sweeping $\alpha$ in practice.
The certificate is conservative for two reasons: (i) it uses the union proxy $\eta(\pi_{\rm LRT})\!\leq\!\alpha T$; (ii) $\Delta_{\hat{Q}}$ can be biased when $\hat{Q}$ overestimates OOD. Thus even when $\alpha>\alpha_{\max}$ and the inequality cannot be certified, LRT may still outperform Q–guidance empirically. In experiments we therefore sweep $\alpha$ on a logarithmic grid and select the knee of the return–risk curve (Fig.  1); the bound serves as a sanity upper bound for $\alpha$, not as a hard constraint.
@@
    $\widehat{\alpha}_{\max}~=~\frac{\widehat{\Delta}_{\hat{Q}}-2\varepsilon_{\rm in}-\kappa\,\widehat{\eta}(\pi_{Q})}{\kappa\,T}\quad$
@@
Why $T$ matters.
For T=50T{=}50 and $\kappa{\approx}1$, even $\Delta_{\hat{Q}}-2\varepsilon_{\rm in}\approx 0.4$ gives $\alpha_{\max}\approx 0.008$ (quite small). Gating only in the last $10$ steps effectively sets T=10T{=}10, yielding $\alpha_{\max}\approx 0.04$. This explains why late–step gating / smaller $\beta_{\max}$ are both stable and easier to certify in practice.
Appendix D Appending with Q-guidance
D.1 Choice of the evaluation center $a_{c}$.
@@ Appendix D Appending with Q-guidance
Appendix E Implementation Details
Training. IQL critic: MLPs (two hidden layers, 256 units), $\gamma\!=\!0.99$, expectile $0.7$ (following IQL [7]), Adam ($3{\times}10^{-5}$), batch 1024, $30$ epochs. Advantages are computed on standardized $(s,a)$. Labels: global top‑$p$ with p=0.2p\!=\!0.2 unless stated. Diffusion: T=50T\!=\!50, DDPM linear schedule on the forward noise $(1-\alpha_{t})$ with endpoints $(1-\alpha_{1}\!=\!10^{-4},\,1-\alpha_{T}\!=\!2{\times}10^{-2})$, , backbone MLP with SiLU activations, two $\epsilon$‑heads; AdamW ($2{\times}10^{-4}$), batch 1024, 150 epochs. Class‑balancing with an EMA of the positive rate; optional soft advantage weight within positives with temperature $\tau_{A}$ and cap $u_{\max}$.
Inference and calibration. We use the posterior variance $\tilde{\sigma}_{t}^{2}$ (DDPM reverse variance) for both heads and LLR; alternative variance parameterizations have also been explored [21]. Gate defaults: $\beta_{\max}\!=\!1$, $\delta\!\in\![1,2]$ (ablate). We calibrate $\hat{\tau}$ on a held‑out state set (disjoint from any rollout episodes) with a fixed‑point update
@@ Appendix E Implementation Details
Evaluation protocol. Unless stated, each configuration is evaluated over $5$ random seeds $\times$ 1010 episodes per seed. We report means $\pm$ standard errors across all episodes. Actions are mapped back to the environment scale and clipped to bounds before stepping. For D4RL environments we also report normalized scores (when supported).

### v2

@@ Appendix C Proofs for Finite-sample Calibration and Stability
C.1. Finite-sample Calibration: Proof of Theorem 5.2.1
@@
C.2. Deterministic Displacement bound: Proof for Lemma 5.2.2
@@
C.3. One-step LLR Decomposition and Moment Bounds
@@
$\Delta\ell_{t}=\underbrace{\frac{\Delta\mu_{t}^{\top}\!\left(m_{t}-\tfrac{\mu_{c,t}+\mu_{u,t}}{2}\right)}{\sigma_{t}^{2}}}_{=:\penalty 10000\ A_{t}\text{ (deterministic given }\mathcal{F}_{t})}\;+\;\underbrace{\frac{\Delta\mu_{t}^{\top}z_{t}}{\sigma_{t}}}_{=:\penalty 10000\ B_{t}\text{ (zero-mean Gaussian)}}.$
@@
where in the last inequality we used the displacement bound $\|m_{t}-\mu_{u,t}\|\leq B_{t}$ from §C.2.
@@
C.4. Sub-Gaussian concentration for the cumulative LLR
@@
Under the scale controls in §C.2,
@@
The cap $\beta_{\max}$ and gradient clipping (via $G,\lambda_{\max}$) yield a per-step deterministic displacement $B_{t}$ and a uniform bound $B_{\text{step}}$; together with the equal-covariance identity, these give (i) explicit bounds on the conditional mean and variance of each LLR increment, and (ii) a sub-Gaussian concentration for the cumulative evidence with proxy $V=\sum_{t}\|\Delta\mu_{t}\|^{2}/\sigma_{t}^{2}$. These stability controls make the calibration reproducible (Theorem 5.2.1) and the gate numerically well-conditioned.
Appendix D When LRT Dominate Q guidance
D.1. Proof for Proposition 5.3
@@ Appendix D When LRT Dominate Q guidance
D.2. Proof of Proposition 5.3
Proof.
Let $\mathcal{E}$ be the event that the generated action is out-of-distribution, i.e., $\mathcal{E}=\{a_{0}\notin\mathcal{S}(s)\}$. Let $\mathcal{R}$ be the event that the trajectory-level likelihood-ratio test rejects the null hypothesis $H_{0}$ (i.e., the gate allows deviation from the background), defined as $\mathcal{R}=\{\ell_{\mathrm{cum}}\geq\tau\}$.
By the Law of Total Probability:
$\eta(\pi_{\rm LRT})=\Pr(\mathcal{E})=\Pr(\mathcal{E}\mid\mathcal{R})\Pr(\mathcal{R})+\Pr(\mathcal{E}\mid\neg\mathcal{R})\Pr(\neg\mathcal{R}).$
  1. (1)
    Rejection Region ($\mathcal{R}$): By the calibration procedure (Algorithm 1 and Theorem 5.2.1), the threshold $\tau$ ensures $\Pr(\mathcal{R})\approx\alpha$. Since $\Pr(\mathcal{E}\mid\mathcal{R})\leq 1$, the first term is bounded by $\alpha$.
  2. (2)
    Acceptance Region ($\neg\mathcal{R}$): The Monotone Support Assumption states that if the test does not reject (gates remain mostly closed), the background head keeps the sample within support, i.e., $\Pr(\mathcal{E}\mid\neg\mathcal{R})\approx 0$.
Combining these, $\eta(\pi_{\rm LRT})\lesssim 1\cdot\alpha+0=\alpha$.
Remark (Union bound): If instead one bounds OOD using per-step Type-I rate $\alpha$ and assumes independence (or uses union bound), the trajectory-level OOD can be conservatively upper-bounded as
$\eta(\pi_{\rm LRT})\leq\sum_{t=1}^{T}\Pr(\text{step }t\text{ false activation})\leq\alpha T.$
This bound is much more conservative than the calibrated $\alpha$ used in practice.
D.3. Condition for LRT to Dominate
[…]
Under Assumption 5.3 with $c(\alpha)\leq\alpha$, a sufficient condition for $\Delta_{\text{true}}\geq 0$ is
$\alpha\penalty 10000\ \leq\penalty 10000\ \alpha_{\max}:=\frac{\Delta_{\widehat{Q}}-2\varepsilon_{\rm in}-\Delta_{\varepsilon}\,\eta(\pi_{Q})}{\Delta_{\varepsilon}\,}.$ (22)
@@ D.3. Condition for LRT to Dominate
Plug $\eta(\pi_{\rm LRT})\leq\alpha$ into (21) and solve $\Delta_{\text{true}}\geq 0$ for $\alpha$. ∎
@@
D.4. A tighter, Data-driven Alternative
The surrogate $\eta(\pi_{\rm LRT})\leq\alpha$ is often loose. Calibrate once and measure $\widehat{\eta}_{\rm LRT}(\alpha)$ on held-out states (our kNN proxy). Then the same derivation yields the data-driven bound
$\Delta_{\text{true}}\penalty 10000\ \geq\penalty 10000\ \Delta_{\widehat{Q}}\penalty 10000\ -\penalty 10000\ 2\varepsilon_{\rm in}\penalty 10000\ -\penalty 10000\ \Delta_{\varepsilon}\big(\widehat{\eta}_{\rm LRT}(\alpha)+\widehat{\eta}(\pi_{Q})\big),$ (23)
@@ D.4. A tighter, Data-driven Alternative
D.5. Optional Refinement with Consistency or Distance-based Bounds
@@
$\Delta_{\text{true}}\penalty 10000\ \gtrsim\penalty 10000\ \Delta_{\widehat{Q}}\penalty 10000\ -\penalty 10000\ 2r_{n}\penalty 10000\ -\penalty 10000\ \Delta_{\varepsilon}\big(\eta(\pi_{\rm LRT})+\eta(\pi_{Q})\big),$
@@
$\Delta_{\text{true}}\penalty 10000\ \geq\penalty 10000\ \Delta_{\widehat{Q}}-2\varepsilon_{\rm in}-L_{Q}\,\mathbb{E}_{s}\!\big[\mathrm{dist}(a_{\rm LRT},\mathcal{S}(s))+\mathrm{dist}(a_{Q},\mathcal{S}(s))\big].$
@@
D.6. Further Discussion for LRT dominate Q
Combining Prop. 5.3 and Prop. 5.3 yields the sufficient condition
$\Delta_{Q^{\rm true}}\penalty 10000\ \geq\penalty 10000\ \Delta_{\hat{Q}}\penalty 10000\ -\penalty 10000\ 2\varepsilon_{\rm in}\penalty 10000\ -\penalty 10000\ \nu\big(\alpha+\eta(\pi_{Q})\big),$
$\alpha_{\max}\penalty 10000\ =\penalty 10000\ \frac{\Delta_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\eta(\pi_{Q})}{\nu},$
@@ D.6. Further Discussion for LRT dominate Q
  • OOD penalty $\nu(\alpha+\eta(\pi_{Q}))$. The LRT part scales with $\alpha$ via the calibrated bound $\eta(\pi_{\rm LRT})\lesssim\alpha$; the Q–guided baseline pays $\nu\,\eta(\pi_{Q})$. Smaller $\nu$—e.g., via a smaller $\beta_{\max}$ or late–step gating when $\sigma_{t}$ is small—makes the condition easier to satisfy.
When does LRT “dominate” Q–guidance?
@@
  • A small–OOD Q baseline helps the certificate. If $\eta(\pi_{Q})$ is small (Q–guidance rarely leaves support), the denominator in $\alpha_{\max}$ is fixed while the numerator increases, so $\alpha_{\max}$ grows: $\eta(\pi_{Q})\!\downarrow\Rightarrow\alpha_{\max}\!\uparrow$. In the extreme $\eta(\pi_{Q}){=}0$, the feasible range reads $\alpha\leq(\Delta_{\hat{Q}}-2\varepsilon_{\rm in})/\nu$. Intuitively, when Q–guidance already behaves conservatively, we can open the LRT gate more (choose a larger $\alpha$) and still retain a certificate.
  • Levers that enlarge $\alpha_{\max}$. (i) Lower $\nu$ (e.g., smaller $\beta_{\max}$ or restricting gating to small–variance steps; anchoring gradients at $\mu_{u}$ also lowers off–support effects); (ii) Increase $\Delta_{\hat{Q}}$ by taking a modest Q–step near $\mu_{\rm LRT}$ when the gate opens or by using evidence–tied/annealed guidance.
Why we still sweep $\alpha$ in practice.
The certificate is conservative for two reasons: (i) it uses the calibrated bound $\eta(\pi_{\rm LRT})\!\lesssim\!\alpha$; (ii) $\Delta_{\hat{Q}}$ can be biased when $\hat{Q}$ overestimates OOD. Thus even when $\alpha>\alpha_{\max}$ and the inequality cannot be certified, LRT may still outperform Q–guidance empirically. In experiments we therefore sweep $\alpha$ on a logarithmic grid and select the knee of the return–risk curve (Fig. 2); the bound serves as a sanity upper bound for $\alpha$, not as a hard constraint.
@@
    $\widehat{\alpha}_{\max}\penalty 10000\ =\penalty 10000\ \frac{\widehat{\Delta}_{\hat{Q}}-2\varepsilon_{\rm in}-\nu\,\widehat{\eta}(\pi_{Q})}{\nu}\quad$
@@
Appendix E Appending with Q-guidance
E.1. Choice of the evaluation center $a_{c}$.
@@ Appendix E Appending with Q-guidance
Appendix F Implementation Details
Training. IQL critic: MLPs (two hidden layers, 256 units), $\gamma\!=\!0.99$, expectile $0.7$ (following IQL Kostrikov2022IQL), Adam ($3{\times}10^{-5}$), batch 1024, $30$ epochs. Advantages are computed on standardized $(s,a)$. Labels: global top‑$p$ with p=0.2p\!=\!0.2 unless stated. Diffusion: T=50T\!=\!50, DDPM linear schedule on the forward noise $(1-\alpha_{t})$ with endpoints $(1-\alpha_{1}\!=\!10^{-4},\,1-\alpha_{T}\!=\!2{\times}10^{-2})$, , backbone MLP with SiLU activations, two $\epsilon$‑heads; AdamW ($2{\times}10^{-4}$), batch 1024, 150 epochs. Class‑balancing with an EMA of the positive rate; optional soft advantage weight within positives with temperature $\tau_{A}$ and cap $u_{\max}$.
Inference and calibration. We use the posterior variance $\tilde{\sigma}_{t}^{2}$ (DDPM reverse variance) for both heads and LLR; alternative variance parameterizations have also been explored (Nichol2021ImprovedDDPM). Gate defaults: $\beta_{\max}\!=\!1$, $\delta\!\in\![1,2]$ (ablate). We calibrate $\hat{\tau}$ on a held‑out state set (disjoint from any rollout episodes) with a fixed‑point update
@@ Appendix F Implementation Details
Evaluation protocol. Unless stated, each configuration is evaluated over $10$ random seeds $\times$ $10$ episodes per seed. We report means $\pm$ standard deviation across all seeds. Actions are mapped back to the environment scale and clipped to bounds before stepping.
Appendix G Additional Experimental Discussion
G.1. Interpreting Table 1.
On all three *-medium-replay-v2 tasks, LRT consistently yields the lowest state-conditional OOD (reported in Table 1 in units of $\times 10^{-2}$), while keeping returns competitive. For Hopper-replay, LRT attains $329\pm 22$ return at $1.84\pm 0.15$ OOD ($\times 10^{-2}$), versus Q at $363\pm 34$ and $6.32\pm 0.53$. For Walker2d-replay, LRT achieves $315\pm 21$ at $0.41\pm 0.07$, whereas Q and LRT+Q raise return to $\approx 3.7\times 10^{2}$ but at $\approx 3.8$ OOD (i.e., $\approx 3.8\times 10^{-2}$). On HalfCheetah-replay, LRT remains conservative in OOD ($1.14\pm 0.25$), while Q and LRT+Q incur much higher OOD ($\approx 13$, i.e., $\approx 13\times 10^{-2}$) for modest return gains.
On the *-medium-v2 tasks, adding a critic step (LRT+Q) provides the strongest return improvements, while LRT alone remains the conservative anchor. For Hopper-medium, LRT+Q reaches the best return ($1197\pm 47$) with higher OOD ($11.32\pm 0.83$) than LRT ($744\pm 71$, $9.21\pm 0.46$). For HalfCheetah-medium, both Q and LRT+Q substantially increase return over LRT (from $3526\pm 38$ to >4400>4400) with a moderate OOD increase (from $3.61\pm 0.97$ to $\approx 5$). For Walker2d-medium, Q and LRT+Q achieve much higher returns ($2282\pm 178$ and $2448\pm 196$) than LRT ($568\pm 49$) while operating at lower OOD ($\approx 5$ vs. $10.79\pm 2.51$; all $\times 10^{-2}$), indicating that value-driven exploitation dominates this dataset.
Overall, LRT serves as a low-risk anchor that suppresses off-support behavior, and LRT+Q moves along the return–risk frontier when critic guidance is beneficial.
G.2. Seed-to-seed Variability on halfcheetah-medium-replay-v2
Why is the std large, and why report relative gaps?
[…]
Dataset LRT vs Q (%) LRT+Q vs Q (%)
hopper-medium-replay-v2 $8.52\pm 12.63$ $3.01\pm 9.33$
halfcheetah-medium-replay-v2 $6.72\pm 22.05$ $2.85\pm 14.67$
walker2d-medium-replay-v2 $16.36\pm 11.79$ $0.25\pm 4.21$
hopper-medium-v2 $36.73\pm 7.56$ $2.78\pm 4.02$
halfcheetah-medium-v2 $19.93\pm 1.12$ $1.51\pm 1.08$
walker2d-medium-v2 $75.11\pm 7.93$ $7.30\pm 7.09$
G.3. Case: walker2d-medium-v2
Interpreting LRT vs. Q on Walker2d-medium.
In Table 1, the reported OOD is in units of $\times 10^{-2}$. For walker2d-medium-v2, the return-optimal setting for LRT may correspond to a relatively large $\alpha$, which increases gate activations and can raise OOD, yet still falls short of the exploitation provided by explicit value gradients. This highlights a key point: LRT controls a calibrated activation risk, while the critic step primarily drives return; when value guidance is reliable and highly beneficial, LRT alone can be under-exploitative, whereas LRT+Q moves along the return–risk frontier.
Appendix H Ablation
H.1. Sensitivity to the “Good Data” Fraction $p$
In our method, $p$ is only used once to mark the top-$p$ transitions as “good” when training the conditional head. We set $p=0.2$ as a simple default that balances dataset coverage while emphasizing high-advantage samples.
We ran an ablation on hopper-medium-replay-v2 with the same training setup. The averaged results are shown in Table 3.
$p$ Return (mean) Return (std) OOD (mean) OOD (std)
0.05 344.1 12.0 0.0172 0.0010
0.10 306.2 11.8 0.0193 0.0029
0.20 342.4 22.9 0.0189 0.0020
0.30 307.1 12.6 0.0168 0.0018
Table 3. Ablation over the “good data” fraction $p$: $p\in\{0.05,0.10,0.20,0.30\}$, $\alpha\in\{0.10,0.05,0.02,0.005\}$
The returns and state-conditional OOD rates are broadly stable across $p$. While $p=0.2$ is slightly better, other choices also perform comparably, indicating that our method is not overly sensitive to the exact choice of $p$.
H.2. Stress Test under Variance Mismatch
We conducted a stress test on hopper-medium-replay-v2: we deliberately violate the equal-variance assumption at sampling time by inflating the background head’s noise standard deviation by a factor $s\in\{1,2,3\}$, while keeping the calibration procedure (which assumes equal variances) unchanged, see Table 4.
As expected, returns drop sharply, Type-I error inflates, and OOD increases as $s$ grows. At scale $s=1$, risk control holds; at $s=2,3$, both Type-I error and OOD increase, leading to large return degradation. This confirms that the guarantee is model-conditional and degrades under deliberate mis-specification. This experiment reflects a stress test rather than a realistic configuration.
scale $\alpha$ $\tau$ Return Type-I OOD
1 0.10 5.5769 361.46 0.0005 0.0202
1 0.05 6.7092 329.43 0.0010 0.0154
1 0.01 9.8676 316.82 0.0000 0.0204
2 0.10 5.5769 127.75 0.0560 0.2023
2 0.05 6.7092 137.83 0.0380 0.1871
2 0.01 9.8676 144.92 0.0085 0.1945
3 0.10 5.5769 70.81 0.1490 0.3362
3 0.05 6.7092 62.72 0.1030 0.3437
3 0.01 9.8676 55.16 0.0455 0.3462
Table 4. Stress test under background variance scaling.